Math 141 (§§4 & 5)  University of South Carolina
D. Meade  Spring 1997

Homework Problems for Chapter 2 Functions and Limits

Section Page Problems
2.1  45 # 3, 4, 9, 10, 11, 14, 16–18, 20, 22, 23, 26, 27, 32, 33, 35–38, 42

Concepts Review: # 1, 2, 4

2.2  52 # 1, 4, 6, 11, 15, 18, 20, 21, 23, 28, 31–33

Concepts Review: # 1–4

2.3  59 # 1, 2, 10abef, 11, 13, 17, 20, 22, 24, 30, 31

Concepts Review: # 1–4

2.4  66 # 1–16, 27–29, 32, 34, 39

Concepts Review: # 1–4

2.5  74 # 7, 10, 13, 18, 20, 21, 23, 29a

Concepts Review: # 1–4

2.6  81 # 2, 3, 6, 12–14, 17–20, 23, 28, 29, 32, 39, 42–44, 50

Concepts Review: # 1–4

2.7  87 # 1–6, 9, 11–13, 15–17, 19, 21–24, 29, 33, 35, 36, 40

Concepts Review: # 1, 3, 4

Review 89 Concepts Test: # 1–9, 11–21, 24–27

Sample Test Problems: # 1–13, 15–25, 29, 29, 311–34

\underline{Notes}

• The assigned problems represent a wide level of difficulty. Many are routine problems. In addition, you will also find i) applications that require you to find and solve an appropriate formulation of the problem, ii) multi-step problems that involve the use of several different ideas and techniques (possibly in new combinations), iii) problems that involve parameters with unspecified values, and iv) problems that require careful understanding (and re-reading) of the general results that have been developed.

• Each chapter concludes with an extended collection of review problems. These problems — both the Concepts Test and the Sample Test Problems — are ideally suited for review before an exam.

• If you need more practice, there are many other problems in the text that are similar to the ones I have assigned. Also, the Mathematics Library (LeConte, third floor) has a collection of Calculus textbooks. These can be a great resource for students needing additional practice.

• Do not hesitate to ask questions – in class, in recitation, by e-mail, during office hours, or in the Math Lab.